# Fine-grained Abnormality Prompt Learning for Zero-shot Anomaly Detection

### v1

@@ 4.1 Main Results
Pixel-level ZSAD Performance. We also compare the pixel-level ZSAD results of our FAPrompt with SotA methods across 14 AD datasets in Table 2. Similar observations can be derived as the image-level results. In particular, CLIP and CLIP-AC are the weakest among the handcrafted text prompt-based methods, primarily due to inappropriate text prompt designs. With better prompt engineering (and adaptation to AD in some cases), WinCLIP, APRIL-GAN, and AnoVL demonstrate better performance. For the learnable text prompt approaches, CoOp shows poor performance due to overfitting on the adaptation dataset, while CoCoOp mitigates this limitation by introducing instance-conditional information, achieving substantial improvement over CoOp and competitive performance to AnomalyCLIP. FAPrompt demonstrates superior performance in identifying a wide range of pixel-level anomalies, significantly outperforming SotA models across nearly all datasets. It surpasses the best competing methods by up to 2.7% AUROC and 4.4% AP on the industrial AD datasets, and by 3.0% AUROC and 3.7% AP on the medical AD datasets. This demonstrates the effectiveness of the fine-grained abnormality prompt in FAPrompt that adaptively capture detailed abnormality semantics in different datasets.
Performance of Learning Complementary Abnormalities. To assess the complementarity of the abnormality prompts learned by FAPrompt, we empirically evaluate the discriminability of each abnormality prompt and its difference to the rest of
Figure 3: t-SNE visualization of prompt-wise anomaly scores on BTAD (01).
other prompts. To this end, we calculate the patch-level anomaly scores $\mathbf{S}_{x}^{a}$ of an image based on the similarity of its patch token embeddings to each individual abnormal prompt embedding (rather than the prototype of the abnormal prompt embeddings) in CAP, and subsequently project the anomaly scores of each sample into a two-dimensional space via t-SNE. As depicted in Fig. 3 on an exemplar dataset, two key observations can be derived: i) despite having slight overlapping, the normal and abnormal samples are distributed into a different group for each individual abnormality prompt, indicating the learning of different abnormal patterns per abnormality prompt; and ii) there is clear separation between normal and abnormal samples for the use of each abnormality prompt in anomaly scoring, indicating the good discriminability of each prompt learned in FAPrompt. Similar patterns can be found in other datasets (see Appendix C.3).
To learn more abnormalities, one straightforward solution is to learn multiple abnormality prompts by using an ensemble of 10 abnormality prompts with each learned by AnomalyCLIP tuned on the auxiliary dataset with a different random seed. We argue that this simple strategy leads to the learning of highly redundant abnormality prompts, rather than the complementary prompts learned in FAPrompt. This is justified by our results of comparing FAPrompt to this ensemble method in Table 4.1. It is clear that the ensemble method can improve AnomalyCLIP to some extent,
Table 3: FAPrompt vs. AnomalyCLIP ensemble.
@@
FAPrompt (88.2, 87.2) (95.0, 85.0) (90.9, 93.0) (85.4, 65.9)
but its abnormality prompts are much less effective than FAPrompt in both image- and pixel-level ZSAD results, despite that no ensemble strategy is used in FAPrompt. This showcases the effectiveness of the abnormality prompts learned in FAPrompt in capturing the fine-grained abnormality details.
4.2 Ablation Study
[…]
Model Industrial Datasets Medical Datasets
Image-level Pixel-level Image-level Pixel-level
Base (85.0, 83.6) (94.4, 84.8) (87.7, 90.6) (83.2, 62.9)
+ CAP (88.1, 87.0) (94.6, 83.9) (90.6, 93.1) (83.8, 63.8)
[…]
+ DAP (86.9, 85.2) (94.8, 84.9) (90.2, 92.3) (84.6, 64.8)
[…]
FAPrompt (88.2, 87.2) (95.0, 85.0) (90.9, 93.0) (85.4, 65.9)
[…]
Hyperparameter Sensitivity Analysis. We analyze the sensitivity of two key hyperparameters of FAPrompt in terms of image-level performance in Fig. 4, including the number of abnormality prompts $K$ in CAP and the number of selected patch tokens $M$ in DAP (similar results can be found for pixel-level results in Appendix C.4). In particular, the performance gets improved with increasing $K$, typically peaking at $K=10$. The performance
Figure 4: Averaged results on industrial and medical datasets with varying $K$ and $M$.
may slightly declines when $K$ is chosen beyond 10. This suggests that while increasing the number of prompts helps capture a wider range of abnormalities, too large $K$ values may introduce noise or redundancy into the prompts. As for the number of selected tokens, $M$, the performance exhibits a similar pattern, with the best performance obtained at a medium value. This suggests that selecting too many abnormal patch candidates may introduce noise or less relevant patches into CAP, leading to the learning of less effective fine-grained anomalities.
Additionally, we also evaluate the sensitivity of the length of learnable normal and abnormal tokens $\{E,E^{\prime}\}$ in CAP module. The Image-level (‘I-AUROC’ and ‘I-AP’)
Figure 5: Averaged results of FAPrompt with varying prompt sizes of $(E,E^{\prime})$.
and pixel-level (‘P-AUROC’ and ‘P-PRO’) ZSAD results are shown in Fig. 5. Overall, the setting of (5, 2) works best for both industrial and medical AD, yielding strong ZSAD performance. Longer prompt lengths, such as (10, 4), can introduce more complexity without clear performance improvement, particularly in pixel-level performance. Using shorter prompt lengths, e.g., the setting of (2, 1), lacks sufficient capacity to support the ZSAD task, leading to consistently weaker performance.

### v2

@@ 4.1 Main Results
Pixel-level ZSAD Performance. We also compare the pixel-level ZSAD results of our FAPrompt with SotA methods across 14 AD datasets in Table 2. Similar observations can be derived as the image-level results. In particular, CLIP and CLIP-AC are the weakest among the handcrafted text prompt-based methods, primarily due to inappropriate text prompt designs. With better prompt engineering (and adaptation to AD in some cases), WinCLIP, APRIL-GAN, and AnoVL demonstrate better performance.
For the learnable text prompt approaches, CoOp shows poor performance due to overfitting on the adaptation dataset, whereas CoCoOp mitigates this by introducing instance-conditional information, achieving substantial improvement over CoOp and competitive performance to BLIP, AnomalyCLIP and FiLO. FAPrompt exhibits superior performance in identifying a wide range of pixel-level anomalies, significantly outperforming SotA models across nearly all datasets, particularly in complex scenarios with larger sample sizes or requiring extensive knowledge transfer, such as DAGM and medical datasets. It surpasses the best competing methods by up to 1.4% AUROC and 4.0% AP on the industrial AD datasets, and by 3.2% AUROC and 4.2% AP on the medical AD datasets. This demonstrates the effectiveness of the fine-grained abnormality prompts in FAPrompt that adaptively capture detailed abnormality semantics in different datasets.
4.2 Analysis of FAPrompt
Module Ablation. We conduct ablation study to assess the contribution of two modules of FAPrompt and report the averaged results across 18 industrial and medical datasets in Table 3. AnomalyCLIP serves as our base model (Baseline), to which we separately add each of our proposed modules: ‘+ CAP’ and ‘+ DAP’.
[…]
Model Industrial Datasets Medical Datasets
Image-level Pixel-level Image-level Pixel-level
Baseline (85.0, 83.6) (94.4, 84.8) (87.7, 90.6) (83.2, 62.9)
+ CAP (88.1, 87.0) (94.6, 83.9) (90.6, 93.1) (83.8, 63.8)
[…]
+ DAP (86.9, 85.2) (94.8, 84.9) (90.2, 92.3) (84.6, 64.8)
[…]
FAPrompt (88.5, 87.5) (95.0, 85.6) (91.0, 93.0) (85.7, 66.2)
[…]
Significance of Sample-wise Abnormality Prior in DAP. To evaluate the effectiveness of our sample-wise abnormality prior in DAP module, we further conduct an ablation study in which we modified the basis of the learned prior by substituting the selected abnormal patch token embeddings (P) with image token embeddings (I), as employed in CoCoOp. The results are presented in Table 4. As shown, CoCoOP yields poor performance in both image- and pixel-level ZSAD tasks. This is primarily due to that the learned prior focuses more on class-level differences while ignoring the sample-specific abnormality details that are essential for effective AD. While the two variants of FAPrompt markedly improve performance over CoCoOP by filtering out normal sample-related noise and introducing fine-grained abnormality semantics through $\mathcal{L}_{prior}$ and CAP, they still fall short the performance of the full FAPrompt.
Model Prior $\mathcal{L}_{prior}$ CAP Industrial Medical
image-level pixel-level image-level pixel-level
CoCoOP I $\times$ $\times$ (73.1, 70.1) (85.5, 80.9) (79.2, 83.5) (79.8, 57.7)
$\text{FAPrompt}^{--}$ I $\times$ (77.3, 75.2) (88.6, 81.0) (83.2, 84.8) (81.6, 62.6)
$\text{FAPrompt}^{-}$ I (87.3, 86.4) (94.2, 83.5) (88.4, 90.3) (83.8, 65.5)
FAPrompt P (88.5, 87.5) (95.0, 85.6) (91.0, 93.0) (85.7, 66.2)
Table 4: Image-level (AUROC, AP) and pixel-level (AUROC, PRO) results using different abnormality prior in DAP.
FAPrompt vs Ensemble methods. To learn more abnormalities, a straightforward solution is to ensemble existing ZSAD methods. Accordingly, we investigate two ensemble strategies in AnomalyCLIP for comparison: i) to learn an ensemble of AnomalyCLIP models with each model tuning an abnormality prompt on the auxiliary dataset using a different random seed (‘AnomalyCLIP Ensemble’), and ii) to learn a AnomalyCLIP model with a set of multiple abnormality prompts constrained to be orthogonal via $\mathcal{L}_{oc}$ (‘AnomalyCLIP Ensemble*’). As shown in Table 5, although both ensemble strategies yield performance gains over the base model AnomalyCLIP, the abnormality prompts derived from these methods are considerably less effective than those produced by FAPrompt. This shortfall is primarily because the ensemble approaches tend to generate highly redundant prompts, whereas FAPrompt learns complementary prompts that capture fine-grained abnormality details unattainable by the ensemble methods.
@@ 4.2 Analysis of FAPrompt
AnomalyCLIP Ensemble* (85.5, 82.6) (94.6, 84.5) (88.8, 91.0) (83.5, 65.6)
Best-matched Prompts (85.1, 84.9) (95.0, 83.2) (85.6, 89.1) (84.7, 65.9)
Weighted Prompts (87.7, 86.7) (94.4, 83.2) (90.9, 92.3) (85.1, 65.7)
FAPrompt (88.5, 87.5) (95.0, 85.6) (91.0, 93.0) (85.7, 66.2)
Table 5: Image-level (AUROC, AP) and pixel-level (AUROC, PRO) results comparison with alternatives.
Prototypical Abnormality Prompt vs Alternative Methods. To verify the effectiveness of averaging the prompts to obtain a unified prototypical abnormality prompt, we compare two alternative variants of FAPrompt in Table 5: i) ‘Best-matched Prompts’, which selects the most similar abnormality prompt per patch/image tokens during anomaly scoring, and ii) ‘Weighted Prompts’, which assign importance weights to the abnormality prompts in the averaged operation based on the top $M$ abnormal patch tokens $\textbf{p}_{x}$. The results indicate that ‘Best-matched Prompts’ performs worse than simple averaging approach, as selecting only a single best-match prompt can lead to the loss of complementary information from other abnormality prompts, limiting the model’s ability to detect the full spectrum of abnormalities. Although ‘Weighted Prompts’ partially mitigates this issue by preserving complementary information through weighted aggregation, its performance still falls short of that achieved by simple averaging. This is due to that the weight learning network can overfit the auxiliary training dataset in the zero-shot setting. Thus, we adopt the straightforward averaging strategy in FAPrompt, as it effectively integrates diverse abnormality cues while preserving their complementary information.
Figure 3: Image-level (Left) and Pixel-level (Right) AUROC results based on different value of $K$ and $M$.
Hyperparameter Sensitivity Analysis. We analyze the sensitivity of two key hyperparameters of FAPrompt on industrial datasets in terms of image- and pixel-level AUROC results in Fig. 3, including the number of abnormality prompts $K$ in CAP and the number of selected patch tokens $M$ in DAP. In particular, the performance gets improved with increasing $K$, typically peaking at $K=10$. The performance may slightly declines when $K$ is chosen beyond 10. This suggests that while increasing the number of prompts helps capture a wider range of abnormalities, too large $K$ values may introduce noise or redundancy into the prompts. As for the number of selected tokens, $M$, the performance exhibits a similar pattern, with the best performance obtained at $M=10$. Beyond this point, a slight decline in performance is observed. This suggests that selecting too many abnormal patch candidates may introduce noise or less relevant patches into learned prior, leading to the learning of less effective fine-grained abnormalities (more comprehensive sensitivity analysis can be found in Appendix C.4).